OpenStudy (anonymous):

SQRT6(3-SQRT6)+8SQRT6

OpenStudy (anonymous):

\[\sqrt{6}\left( 3-\sqrt{6} \right)+8\sqrt{6}\]

OpenStudy (anonymous):

\[3\sqrt{6}+8\sqrt{6}-6\]

OpenStudy (anonymous):

\[11\sqrt{6}-6\]

OpenStudy (anonymous):

\[6/X=7/X+3\]

OpenStudy (anonymous):

Multiply by x \[6=7+3x\] Solve this for x.